OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

@dmezzullo Help please.

OpenStudy (anonymous):

@precal

OpenStudy (precal):

use the definition of a polar

OpenStudy (anonymous):



OpenStudy (anonymous):

you can use for the angle \(-\frac{\pi}{4}\) or \(\frac{7\pi}{4}\)

OpenStudy (anonymous):

for \(r\) use \(r=\sqrt{2^2+2^2}\)

OpenStudy (anonymous):

thanks !

OpenStudy (precal):

follow satellite73's lead; you are in good hands